# The algebraic and geometric classification of

### v1

@@ 1.3.14. Compatible pre-Lie algebras defined on $\mathcal{C}_{08}$
    {$e_{1}\cdot e_{1}=e_{1},$&$e_{1}\cdot e_{2}=2e_{2},$&$e_{2}\cdot e_{1}=\frac{1}{2}e_{1}+e_{2},$&$e_{2}\cdot e_{2}=e_{2},$\\ $e_{1}\ast e_{1}=\alpha_{1}e_{1}+\alpha_{2}e_{2},$&$e_{1}\ast e_{2}=2\alpha_{1}e_{2},$&$e_{2}\ast e_{1}=\alpha_{3}e_{1}+\alpha_{1}e_{2},$&$e_{2}\ast e_{2}=2\alpha_{3}e_{2}.$\end{tabular}\right.$
@@
    {$e_{1}\cdot e_{1}=e_{1},$&$e_{1}\cdot e_{2}=2e_{2},$\\ $e_{2}\cdot e_{1}=\frac{1}{2}e_{1}+e_{2},$&$e_{2}\cdot e_{2}=e_{2},$\\ $e_{1}\ast e_{1}=\alpha_{4}e_{1},$&$e_{1}\ast e_{2}=\alpha_{5}e_{1}+2\alpha_{4}e_{2},$\\ $e_{2}\ast e_{1}=\left(2\alpha_{5}+\frac{1}{2}\alpha_{4}\right)e_{1}+\alpha_{4}e_{2},$&$e_{2}\ast e_{2}=\frac{1}{2}\alpha_{5}e_{1}+\left(\alpha_{4}+\alpha_{5}\right)e_{2}.$\end{tabular}\right.$
@@
Moreover, the algebras $\mathcal{C}_{40}^{\alpha,\beta,\gamma}$ and $\mathcal{C}_{40}^{\alpha^{\prime},\beta^{\prime},\gamma^{\prime}}$ are isomorphic if and only if
$\left(\alpha^{\prime},\beta^{\prime},\gamma^{\prime}\right)=\left(\alpha,\beta,\gamma\right)$ or $\left(\alpha^{\prime},\beta^{\prime},\gamma^{\prime}\right)=\left(4\gamma-\alpha,\beta-8\alpha+16\gamma,\gamma\right)$.
Also, the algebras $\mathcal{C}_{41}^{\alpha,\beta}$ and $\mathcal{C}_{41}^{\alpha^{\prime},\beta^{\prime}}$ are isomorphic if and only if
$\left(\alpha^{\prime},\beta^{\prime}\right)=\left(\alpha,\beta\right)$ or $\left(\alpha^{\prime},\beta^{\prime}\right)=\left(\alpha+4\beta,-\beta\right)$.
1.4. The algebraic classification of compatible commutative associative algebras
It is easy to see that each commutative pre-Lie algebra is associative. Hence, we can choose only commutative compatible pre-Lie algebras from the Theorem 7.
Theorem 8.
@@ 1.4. The algebraic classification of compatible commutative associative algebras
$\mathcal{C}_{03}$ :::: $e_{1}\ast e_{1}=e_{2}.$
$\mathcal{C}_{05}^{0}$ :::: $e_{1}\ast e_{1}=e_{1}.$
$\mathcal{C}_{06}^{1}$ :::: $e_{1}\ast e_{1}=e_{1},\ e_{1}\ast e_{2}=e_{2},\ e_{2}\ast e_{1}=e_{2}.$
$\mathcal{C}_{07}$ :::: $e_{1}\ast e_{1}=e_{1},\ e_{2}\ast e_{2}=e_{2}.$
$\mathcal{C}_{13}^{1}$ :::: {$e_{1}\cdot e_{1}=e_{2},$&&\\ $e_{1}\ast e_{1}=e_{1},$&$e_{1}\ast e_{2}=e_{2},$&$e_{2}\ast e_{1}=e_{2}.$\end{tabular}\right.$
$\mathcal{C}_{15}^{\alpha}$ :::: {$e_{1}\cdot e_{1}=e_{2},$\\ $e_{1}\ast e_{1}=\alpha e_{2}.$\end{tabular}\right.$
$\mathcal{C}_{16}^{0}$ :::: {$e_{1}\cdot e_{1}=e_{2},$&\\ $e_{1}\ast e_{1}=e_{1}.$\end{tabular}\right.$
$\mathcal{C}_{18}^{\alpha}$ :::: {$e_{1}\cdot e_{1}=e_{2},$&&&\\ $e_{1}\ast e_{1}=\alpha e_{2},$&$e_{1}\ast e_{2}=e_{1},$&$e_{2}\ast e_{1}=e_{1},$&$e_{2}\ast e_{2}=e_{2}.$\end{tabular}\right.$
$\mathcal{C}_{24}^{0,\beta,0}$ :::: {$e_{1}\cdot e_{1}=e_{1},$\\ $e_{1}\ast e_{1}=\beta e_{1}+e_{2}.$\end{tabular}\right.$
$\mathcal{C}_{25}^{0,\beta,0}$ :::: {$e_{1}\cdot e_{1}=e_{1},$\\ $e_{1}\ast e_{1}=\beta e_{1}.$\end{tabular}\right.$
$\mathcal{C}_{29}^{\alpha}$ :::: {$e_{1}\cdot e_{1}=e_{1},$&\\ $e_{1}\ast e_{1}=\alpha e_{1},$&$e_{2}\ast e_{2}=e_{2}.$\end{tabular}\right.$
$\mathcal{C}_{30}^{\alpha,\beta}$ :::: {$e_{1}\cdot e_{1}=e_{1},$&&&\\ $e_{1}\ast e_{1}=\alpha e_{1}+\beta e_{2},$&$e_{1}\ast e_{2}=e_{1},$&$e_{2}\ast e_{1}=e_{1},$&$e_{2}\ast e_{2}=e_{2}.$\end{tabular}\right.$
$\mathcal{C}_{31}^{1,\beta,\beta}$ :::: {$e_{1}\cdot e_{1}=e_{1},$&$e_{1}\cdot e_{2}=e_{2},$&$e_{2}\cdot e_{1}=e_{2},$\\ $e_{1}\ast e_{1}=\beta e_{1}+e_{2},$&$e_{1}\ast e_{2}=\beta e_{2},$&$e_{2}\ast e_{1}=\beta e_{2}.$\end{tabular}\right.$
$\mathcal{C}_{32}^{1,\beta,\beta}$ :::: {$e_{1}\cdot e_{1}=e_{1},$&$e_{1}\cdot e_{2}=e_{2},$&$e_{2}\cdot e_{1}=e_{2},$\\ $e_{1}\ast e_{1}=\beta e_{1},$&$e_{1}\ast e_{2}=\beta e_{2},$&$e_{2}\ast e_{1}=\beta e_{2}.$\end{tabular}\right.$
$\mathcal{C}_{34}^{\alpha,\beta}$ :::: {$e_{1}\cdot e_{1}=e_{1},$&$e_{1}\cdot e_{2}=e_{2},$&$e_{2}\cdot e_{1}=e_{2},$&\\ $e_{1}\ast e_{1}=\alpha e_{1},$&$e_{1}\ast e_{2}=\alpha e_{2},$&$e_{2}\ast e_{1}=\alpha e_{2},$&$e_{2}\ast e_{2}=e_{1}+\beta e_{2}.$\end{tabular}\right.$
$\mathcal{C}_{35}^{\alpha}$ :::: {$e_{1}\cdot e_{1}=e_{1},$&$e_{1}\cdot e_{2}=e_{2},$&$e_{2}\cdot e_{1}=e_{2},$&\\ $e_{1}\ast e_{1}=\alpha e_{1},$&$e_{1}\ast e_{2}=\alpha e_{2},$&$e_{2}\ast e_{1}=\alpha e_{2},$&$e_{2}\ast e_{2}=e_{2}.$\end{tabular}\right.$
$\mathcal{C}_{38}^{\alpha,\beta,\gamma}$ :::: {$e_{1}\cdot e_{1}=e_{1},$&$e_{2}\cdot e_{2}=e_{2},$&&\\ $e_{1}\ast e_{1}=\left(\gamma+\beta-\alpha\right)e_{1}-\beta e_{2},$&$e_{1}\ast e_{2}=\alpha e_{1}+\beta e_{2},$\\ $e_{2}\ast e_{1}=\alpha e_{1}+\beta e_{2},$&$e_{2}\ast e_{2}=-\alpha e_{1}+\gamma e_{2}.$\end{tabular}\right.$
$\mathcal{C}_{39}^{\alpha,\beta\neq\alpha}$ :::: {$e_{1}\cdot e_{1}=e_{1},$&$e_{2}\cdot e_{2}=e_{2},$\\ $e_{1}\ast e_{1}=\alpha e_{1},$&$e_{2}\ast e_{2}=\beta e_{2}.$\end{tabular}\right.$
@@
1.5. The algebraic classification of compatible associative algebras
Definition 9.
A compatible associative algebra is a vector space ${\bf A}$ equipped with two multiplication $\cdot$ and another multiplication $\ast$,∗ , such that, $({\bf A},\cdot),$ $({\bf A},\ast)$ and $({\bf A},\cdot+\ast)$ are associative algebras. These two operations are required to satisfy the following identities.
$\left(x\cdot y\right)\cdot z$ $=$ $x\cdot\left(y\cdot z\right),$
$\left(x\ast y\right)\ast z$ $=$ $x\ast\left(y\ast z\right),$
@@ 1.5. The algebraic classification of compatible associative algebras
Theorem 10.
Let $\mathcal{C}$ be a nonzero $2$-dimensional compatible associative algebra. Then $\mathcal{C}$ is isomorphic to one and only one compatible commutative associative algebra listed in Theorem 8 or one of the following algebras:
$\mathcal{C}_{05}^{1}$ :::: $e_{1}\ast e_{1}=e_{1},\ e_{1}\ast e_{2}=e_{2}.$
$\mathcal{C}_{06}^{0}$ :::: $e_{1}\ast e_{1}=e_{1},\ \ e_{2}\ast e_{1}=e_{2}.$
$\mathcal{C}_{25}^{1,\beta,\beta}$ :::: {$e_{1}\cdot e_{1}=e_{1},$&$e_{1}\cdot e_{2}=e_{2},$\\ $e_{1}\ast e_{1}=\beta e_{1},$&$e_{1}\ast e_{2}=\beta e_{2}.$\end{tabular}\right.$
$\mathcal{C}_{28}$ :::: {$e_{1}\cdot e_{1}=e_{1},$&$e_{1}\cdot e_{2}=e_{2},$\\ $e_{2}\ast e_{1}=e_{1},$&$e_{2}\ast e_{2}=e_{2}.$\end{tabular}\right.$
$\mathcal{C}_{32}^{0,\beta,0}$ :::: {$e_{1}\cdot e_{1}=e_{1},$&$e_{2}\cdot e_{1}=e_{2},$\\ $e_{1}\ast e_{1}=\beta e_{1},$&$e_{2}\ast e_{1}=\beta e_{2}.$\end{tabular}\right.$
$\mathcal{C}_{33}$ :::: {$e_{1}\cdot e_{1}=e_{1},$&$e_{2}\cdot e_{1}=e_{2},$\\ $e_{1}\ast e_{2}=e_{1},$&$e_{2}\ast e_{2}=e_{2}.$\end{tabular}\right.$
1.6. The algebraic classification of compatible Novikov algebras
Definition 11.
A compatible Novikov algebra is a vector space ${\bf A}$ equipped with two multiplication $\cdot$ and another multiplication $\ast$,∗ , such that, $({\bf A},\cdot),$ $({\bf A},\ast)$ and $({\bf A},\cdot+\ast)$ are Novikov algebras. These two operations are required to satisfy the following identities.
$\left(x\cdot y\right)\cdot z-x\cdot\left(y\cdot z\right)$ $=$ $\left(y\cdot x\right)\cdot z-y\cdot\left(x\cdot z\right),$
$(x\cdot y)\cdot z$ $=$ $(x\cdot z)\cdot y,$
$\left(x\ast y\right)\ast z-x\ast\left(y\ast z\right)$ $=$ $\left(y\ast x\right)\ast z-y\ast\left(x\ast z\right),$
$(x\ast y)\ast z$ $=$ $(x\ast z)\ast y,$
$(x\ast y)\cdot z-x\ast(y\cdot z)+(x\cdot y)\ast z-x\cdot(y\ast z)$ $=$ $(y\ast x)\cdot z-y\ast(x\cdot z)+(y\cdot x)\ast z-y\cdot(x\ast z),$

### v2

@@ 2.3.14 Compatible pre-Lie algebras defined on $\boldsymbol{\mathcal{C}_{08}}$
    $e_{1}\cdot e_{1}=e_{1}$, $e_{1}\cdot e_{2}=2e_{2}$, $e_{2}\cdot e_{1}=\frac{1}{2}e_{1}+e_{2}$, $e_{2}\cdot e_{2}=e_{2}$, $e_{1}\ast e_{1}=\alpha_{1}e_{1}+\alpha_{2}e_{2}$, $e_{1}\ast e_{2}=2\alpha_{1}e_{2}$, $e_{2}\ast e_{1}=\alpha_{3}e_{1}+\alpha_{1}e_{2}$, $e_{2}\ast e_{2}=2\alpha_{3}e_{2}$,
@@
    $e_{1}\cdot e_{1}=e_{1}$, $e_{1}\cdot e_{2}=2e_{2}$, $e_{2}\cdot e_{1}=\frac{1}{2}e_{1}+e_{2}$, $e_{2}\cdot e_{2}=e_{2}$, $e_{1}\ast e_{1}=\alpha_{4}e_{1}$, $e_{1}\ast e_{2}=\alpha_{5}e_{1}+2\alpha_{4}e_{2}$, $e_{2}\ast e_{1}=\bigl{(}2\alpha_{5}+\frac{1}{2}\alpha_{4}\bigr{)}e_{1}+\alpha_{4}e_{2}$, $e_{2}\ast e_{2}=\frac{1}{2}\alpha_{5}e_{1}+(\alpha_{4}+\alpha_{5})e_{2}$.
@@
$\mathcal{C}_{40}^{\alpha,\beta,\gamma}\colon\ e_{1}\cdot e_{1}=e_{1},\quad e_{1}\cdot e_{2}=2e_{2},\quad e_{2}\cdot e_{1}=\tfrac{1}{2}e_{1}+e_{2},\quad e_{2}\cdot e_{2}=e_{2},$
$\ e_{1}\ast e_{1}=\alpha e_{1}+\beta e_{2},\quad e_{1}\ast e_{2}=2\alpha e_{2},\quad e_{2}\ast e_{1}=\gamma e_{1}+\alpha e_{2},\quad e_{2}\ast e_{2}=2\gamma e_{2},$
$\mathcal{C}_{41}^{\alpha,\beta\neq 0}\colon\ e_{1}\cdot e_{1}=e_{1},\quad e_{1}\cdot e_{2}=2e_{2},\quad e_{2}\cdot e_{1}=\tfrac{1}{2}e_{1}+e_{2},\quad e_{2}\cdot e_{2}=e_{2},$
@@
Moreover, the algebras $\mathcal{C}_{40}^{\alpha,\beta,\gamma}$ and $\mathcal{C}_{40}^{\alpha^{\prime},\beta^{\prime},\gamma^{\prime}}$ are isomorphic if and only if $\bigl{(}\alpha^{\prime},\beta^{\prime},\gamma^{\prime}\bigr{)}=(\alpha,\beta,\gamma)$ or ${\bigl{(}\alpha^{\prime},\beta^{\prime},\gamma^{\prime}\bigr{)}=(4\gamma-\alpha,\beta-8\alpha+16\gamma,\gamma)}$. Also, the algebras $\mathcal{C}_{41}^{\alpha,\beta}$ and $\mathcal{C}_{41}^{\alpha^{\prime},\beta^{\prime}}$ are isomorphic if and only if $\bigl{(}\alpha^{\prime},\beta^{\prime}\bigr{)}=(\alpha,\beta)$ or $\bigl{(}\alpha^{\prime},\beta^{\prime}\bigr{)}=(\alpha+4\beta,-\beta)$.
2.4 The algebraic classification of compatible commutative
associative algebras
It is easy to see that each commutative pre-Lie algebra is associative. Hence, we can choose only commutative compatible pre-Lie algebras from Theorem 2.7.
Theorem 2.8.
@@ 2.4 The algebraic classification of compatible commutative
$\mathcal{C}_{03}\colon\ e_{1}\ast e_{1}=e_{2},$
$\mathcal{C}_{05}^{0}\colon\ e_{1}\ast e_{1}=e_{1},$
$\mathcal{C}_{06}^{1}\colon\ e_{1}\ast e_{1}=e_{1},\quad e_{1}\ast e_{2}=e_{2},\quad e_{2}\ast e_{1}=e_{2},$
$\mathcal{C}_{07}\colon\ e_{1}\ast e_{1}=e_{1},\quad e_{2}\ast e_{2}=e_{2},$
$\mathcal{C}_{13}^{1}\colon\ e_{1}\cdot e_{1}=e_{2},\quad e_{1}\ast e_{1}=e_{1},\quad e_{1}\ast e_{2}=e_{2},\quad e_{2}\ast e_{1}=e_{2},$
$\mathcal{C}_{15}^{\alpha}\colon\ e_{1}\cdot e_{1}=e_{2},\quad e_{1}\ast e_{1}=\alpha e_{2},$
$\mathcal{C}_{16}^{0}\colon\ e_{1}\cdot e_{1}=e_{2},\quad e_{1}\ast e_{1}=e_{1},$
$\mathcal{C}_{18}^{\alpha}\colon\ e_{1}\cdot e_{1}=e_{2},\quad e_{1}\ast e_{1}=\alpha e_{2},\quad e_{1}\ast e_{2}=e_{1},\quad e_{2}\ast e_{1}=e_{1},\quad e_{2}\ast e_{2}=e_{2},$
$\mathcal{C}_{24}^{0,\beta,0}\colon\ e_{1}\cdot e_{1}=e_{1},\quad e_{1}\ast e_{1}=\beta e_{1}+e_{2},$
$\mathcal{C}_{25}^{0,\beta,0}\colon\ e_{1}\cdot e_{1}=e_{1},\quad e_{1}\ast e_{1}=\beta e_{1},$
$\mathcal{C}_{29}^{\alpha}\colon\ e_{1}\cdot e_{1}=e_{1},\quad e_{1}\ast e_{1}=\alpha e_{1},\quad e_{2}\ast e_{2}=e_{2},$
$\mathcal{C}_{30}^{\alpha,\beta}\colon\ e_{1}\cdot e_{1}=e_{1},\quad e_{1}\ast e_{1}=\alpha e_{1}+\beta e_{2},\quad e_{1}\ast e_{2}=e_{1},\quad e_{2}\ast e_{1}=e_{1},\quad e_{2}\ast e_{2}=e_{2},$
$\mathcal{C}_{31}^{1,\beta,\beta}\colon\ e_{1}\cdot e_{1}=e_{1},\quad e_{1}\cdot e_{2}=e_{2},\quad e_{2}\cdot e_{1}=e_{2},$
$\ e_{1}\ast e_{1}=\beta e_{1}+e_{2},\quad e_{1}\ast e_{2}=\beta e_{2},\quad e_{2}\ast e_{1}=\beta e_{2},$
$\mathcal{C}_{32}^{1,\beta,\beta}\colon\ e_{1}\cdot e_{1}=e_{1},\quad e_{1}\cdot e_{2}=e_{2},\quad e_{2}\cdot e_{1}=e_{2},$
$\ e_{1}\ast e_{1}=\beta e_{1},\quad e_{1}\ast e_{2}=\beta e_{2},\quad e_{2}\ast e_{1}=\beta e_{2},$
$\mathcal{C}_{34}^{\alpha,\beta}\colon\ e_{1}\cdot e_{1}=e_{1},\quad e_{1}\cdot e_{2}=e_{2},\quad e_{2}\cdot e_{1}=e_{2},$
$\ e_{1}\ast e_{1}=\alpha e_{1},\quad e_{1}\ast e_{2}=\alpha e_{2},\quad e_{2}\ast e_{1}=\alpha e_{2},\quad e_{2}\ast e_{2}=e_{1}+\beta e_{2},$
$\mathcal{C}_{35}^{\alpha}\colon\ e_{1}\cdot e_{1}=e_{1},\quad e_{1}\cdot e_{2}=e_{2},\quad e_{2}\cdot e_{1}=e_{2},$
$\ e_{1}\ast e_{1}=\alpha e_{1},\quad e_{1}\ast e_{2}=\alpha e_{2},\quad e_{2}\ast e_{1}=\alpha e_{2},\quad e_{2}\ast e_{2}=e_{2},$
$\mathcal{C}_{38}^{\alpha,\beta,\gamma}\colon\ e_{1}\cdot e_{1}=e_{1},\quad e_{2}\cdot e_{2}=e_{2},\quad e_{1}\ast e_{1}=(\gamma+\beta-\alpha)e_{1}-\beta e_{2},$
$\ e_{1}\ast e_{2}=\alpha e_{1}+\beta e_{2},\quad e_{2}\ast e_{1}=\alpha e_{1}+\beta e_{2},\quad e_{2}\ast e_{2}=-\alpha e_{1}+\gamma e_{2},$
$\mathcal{C}_{39}^{\alpha,\beta\neq\alpha}\colon\ e_{1}\cdot e_{1}=e_{1},\quad e_{2}\cdot e_{2}=e_{2},\quad e_{1}\ast e_{1}=\alpha e_{1},\quad e_{2}\ast e_{2}=\beta e_{2}.$
@@
2.5 The algebraic classification of compatible associative algebras
Definition 2.9.
A compatible associative algebra is a vector space ${\bf A}$ equipped with two multiplications: $\cdot$ and another multiplication $\ast$, such that, $({\bf A},\cdot)$, $({\bf A},\ast)$ and $({\bf A},\cdot+\ast)$ are associative algebras. These two operations are required to satisfy the following identities:
$(x\cdot y)\cdot z=x\cdot(y\cdot z),\quad(x\ast y)\ast z=x\ast(y\ast z),$
@@ 2.5 The algebraic classification of compatible associative algebras
Theorem 2.10.
Let $\mathcal{C}$ be a nonzero $2$-dimensional compatible associative algebra. Then $\mathcal{C}$ is isomorphic to one and only one compatible commutative associative algebra listed in Theorem 2.8 or one of the following algebras:
$\mathcal{C}_{05}^{1}\colon\ e_{1}\ast e_{1}=e_{1},\quad e_{1}\ast e_{2}=e_{2},$
$\mathcal{C}_{06}^{0}\colon\ e_{1}\ast e_{1}=e_{1},\quad e_{2}\ast e_{1}=e_{2},$
$\mathcal{C}_{25}^{1,\beta,\beta}\colon\ e_{1}\cdot e_{1}=e_{1},\quad e_{1}\cdot e_{2}=e_{2},\quad e_{1}\ast e_{1}=\beta e_{1},\quad e_{1}\ast e_{2}=\beta e_{2},$
$\mathcal{C}_{28}\colon\ e_{1}\cdot e_{1}=e_{1},\quad e_{1}\cdot e_{2}=e_{2},\quad e_{2}\ast e_{1}=e_{1},\quad e_{2}\ast e_{2}=e_{2},$
$\mathcal{C}_{32}^{0,\beta,0}\colon\ e_{1}\cdot e_{1}=e_{1},\quad e_{2}\cdot e_{1}=e_{2},\quad e_{1}\ast e_{1}=\beta e_{1},\quad e_{2}\ast e_{1}=\beta e_{2},$
$\mathcal{C}_{33}\colon\ e_{1}\cdot e_{1}=e_{1},\quad e_{2}\cdot e_{1}=e_{2},\quad e_{1}\ast e_{2}=e_{1},\quad e_{2}\ast e_{2}=e_{2}.$
2.6 The algebraic classification of compatible Novikov algebras
Definition 2.11.
A compatible Novikov algebra is a vector space ${\bf A}$ equipped with two multiplications: $\cdot$ and another multiplication $\ast$, such that, $({\bf A},\cdot)$, $({\bf A},\ast)$ and $({\bf A},\cdot+\ast)$ are Novikov algebras. These two operations are required to satisfy the following identities:
$(x\cdot y)\cdot z-x\cdot(y\cdot z)=(y\cdot x)\cdot z-y\cdot(x\cdot z),\quad(x\cdot y)\cdot z=(x\cdot z)\cdot y,$
$(x\ast y)\ast z-x\ast(y\ast z)=(y\ast x)\ast z-y\ast(x\ast z),\quad(x\ast y)\ast z=(x\ast z)\ast y,$
$(x\ast y)\cdot z-x\ast(y\cdot z)+(x\cdot y)\ast z-x\cdot(y\ast z)$
$\qquad=(y\ast x)\cdot z-y\ast(x\cdot z)+(y\cdot x)\ast z-y\cdot(x\ast z),$